Show pageDiscussionOld revisionsBacklinksBack to top This page is read only. You can view the source, but not change it. Ask your administrator if you think this is wrong. ======20241206====== =====Time===== =====Attendance===== Vianne, Sarah, Dylan, Krishna =====Spaces Occupied===== control room =====Goals===== =====Equipment Used===== ^ Scope ^ Camera ^ Focus ^ |Main|asi 2600|| |Finder|ASI 482|2465| =====Frame Table===== ^ Time Observed ^ Source Name ^ Filter ^ Int Time [s] ^ Gain ^ Airmass ^ Mag (filter) ^ RA [h] ^ DEC [deg] ^ Source Type ^Scope ^ Purpose ^ Notes^ | YYYYmmddTHHMM | Object | Filter | time | gain | airmass | mag(V) | HH | DD | type | Scope | Purpose | tracking? Others?| | 202412061838 | HIP_99887 | V | 40s | gain | airmass | mag(V) | HH | DD | | Scope | Variable Stars Study | tracking? Others?| | 202412061853 | HIP_104002 | V | 40s | gain | airmass | mag(V) | HH | DD | type | Scope | Variable Stars Study | tracking? Others?| | 202412061906 | HD_203156 | V | 2.5s | gain | airmass | mag(V) | HH | DD | type | Scope | Variable Stars Study | tracking? Others?| | 202412061921 | HIP_105369 | V | 40s | gain | airmass | mag(V) | HH | DD | type | Scope | Variable Stars Study | tracking? Others?| | 202412061950 | HIP_99887 | V | 40s | gain | airmass | mag(V) | HH | DD | | Scope | Variable Stars Study | tracking? Others?| | 202412061958 | HIP_104002 | V | 40s | gain | airmass | mag(V) | HH | DD | type | Scope | Variable Stars Study | tracking? Others?| | 202412062014 | HIP_105369 | V | 40s | gain | airmass | mag(V) | HH | DD | type | Scope | Variable Stars Study | tracking? Others?| | 202412062025 | HD_203156 | V | 2.5s | gain | airmass | mag(V) | HH | DD | type | Scope | Variable Stars Study | tracking? Others?| =====Log===== \\ - Started set up, only one of the mirrors doors opened the west one \\ - Mirror doors were fixed by waiting an hour and trying again after opening and closing a few times \\ - the blog time was originally off, so could not make a blog post, this was fixed \\ - Started variable plan \\ - Had trouble finding the first of the frames, but by reinitializing the telescope and moving it around a little bit we were able to reinitialize it to the right variable star \\ - Started taking darks, flats, darkflats \\ - North and west mirror doors refused to close at the end of the night. Temperature was 30 degrees. \\ - The west door began working again after a few minutes. The north one still isn't working. \\ - All of the doors worked at the end of the night. Seems like letting the dome warm for a while fixed it. =====ToDo===== =====Notes===== =====Weather===== =====Data===== {{tag>[picsarr]}}